OpenStudy (anonymous):

Are the complex fractions (x^2 - x - 20/4) / (x-5/10) and (10/4) / (x-5/x^2 - x - 20) equivalent? Simplify each, and then explain why or why not

OpenStudy (anonymous):

1, 2, 3, and to the 4. snoop doggy dogg and dr dre is at yo door.

OpenStudy (anonymous):

@koreanhugg If you'd do some manipulations, you'd find that the fractions are actually similar. How to do simplification of a fraction: \[a / (b / c) = a . (c / b)\] *notice that the / is changed to a . and positions of b and c are reversed

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

Yep the fractions are similar if you'd do some manipulations you'll see

OpenStudy (anonymous):

I tried but I still don't get it !

OpenStudy (anonymous):

(x^2 - x - 20/4) / (x-5/10) = (x^2 - x - 20 / 4) x (10 / x-5) (10/4) / (x-5/x^2 - x - 20) = (10/4) x (x^2 - x - 20 / x-5) ---- (1) Rearranging (1): = (x^2 - x - 20 / 4) x (10 / x-5) Thus they are equal

OpenStudy (anonymous):

wait so why are they equal again ?

OpenStudy (anonymous):

@dauspex

OpenStudy (anonymous):

Not sure if theres a reason for equality but its just because when you manipulate the expressions you get the same equation?